# The geodesic cover problem for butterfly networks

### v2

@@ Abstract
a Department of Information Science, College of Computing Science and Engineering, Kuwait University, Kuwait
pauldmanuel@gmail.com, p.manuel@ku.edu.kw
b Faculty of Mathematics and Physics, University of Ljubljana, Slovenia
sandi.klavzar@fmf.uni-lj.si
c Faculty of Natural Sciences and Mathematics, University of Maribor, Slovenia
d Institute of Mathematics, Physics and Mechanics, Ljubljana, Slovenia
e Department of Mathematics, Ethiraj College for Women, Chennai, Tamilnadu, India
prabha75@gmail.com
f Department of Mathematics, School of Science and Humanities, Shiv Nadar University Chennai, Kalavakkam, India
andrew23610032@snuchennai.edu.in
@@ 2 Preliminaries
    If $S^{\prime}\subseteq S^{\prime\prime}\subseteq V(G)$ and $Y\subseteq Z(G)$, then ${\rm gcover}(Y,S^{\prime\prime},G)$geq ${\rm gcover}(Y,S^{\prime},G)$.
@@
    ${\rm gcover}(G)$ $=$ ${\rm gcover}(M(G),G)$.
Proof.
Assertions (i) and (ii) are straightforward, hence we consider only (iii). Since $M(G)\subseteq Z(G)$, (ii) implies ${\rm gcover}(G)={\rm gcover}(Z(G),V,G)\leq{\rm gcover}(M(G),V,G)={\rm gcover}(M(G),G)$. Since each geodesic is a subpath of some maximal geodesic, for each geodesic cover $S$ of $Z(G)$, there exists a geodesic cover $S^{\prime}$ of $M(G)$ such that $|S|=|S^{\prime}|$. Therefore, ${\rm gcover}(Z(G),V,G)\geq{\rm gcover}(M(G),V,G)$ and consecutively ${\rm gcover}(G)={\rm gcover}(Z(G),V,G)\geq{\rm gcover}(M(G),V,G)={\rm gcover}(M(G),G)$. ∎
Proposition 2.2.
@@
Proof.
Clearly, each maximal geodesic of $K_{r,r}$ is a (diametral) path of length $2$. Therefore, ${\rm gcover}(K_{r,r})\geq\lceil(2/3)r\rceil$. On the other hand, it is a simple exercise to construct a geodesic cover of cardinality $\lceil(2/3)r\rceil$. ∎
Butterfly is considered as one of the best parallel architectures [6, 7, 15]. For $r\geq 3$, the $r$-dimensional butterfly network $BF(r)$ has vertices $[j,s]$, where $s\in\{0,1\}^{r}$ and $j\in\{0,1,\ldots,r\}$. The vertices $[j,s]$ and $[j^{\prime},s^{\prime}]$ are adjacent if $j^{\prime}=j+1$, and either $s=s^{\prime}$ or $s$ and $s^{\prime}$ differ precisely in the $j^{\rm th}$ bit. $BF(r)$ has $(r+1)2^{r}$ vertices and $r2^{r+1}$ edges. A vertex $[j,s]$ is at level $j$ and row $s$. There are two standard graphical representations for ${\rm BF}(r)$, normal representation and diamond representation, see Fig. 1.
@@
Lemma 2.3.
A geodesic of ${\rm BF}(r)$ contains at most two vertices of level 00 and at most two vertices of level $r$. Moreover, if a geodesic contains two vertices of level 00, then they are the ends of the geodesic (and similarly for level r).
Proof.
Let us assume that there exists a geodesic $P$ which contains more than two vertices of level 00, say $v_{i},v_{j}$, and $v_{k}$. See Fig. 2(b). Then one of these three vertices must be an internal vertex of $P$, say $v_{j}$. The deletion of the vertices at level 00 from ${\rm BF}(r)$ disconnects ${\rm BF}(r)$ into two vertex disjoint components $G_{1}$ and $G_{2}$, where both $G_{1}$ and $G_{2}$ are isomorphic to ${\rm BF}(r-1)$, cf. [7, 13, 16]. Since $v_{j}$ is an internal vertex of $P$ and of degree $2$, its neighbors $v_{j-1}$ and $v_{j+1}$ also lie in $P$. Moreover, one of the adjacent vertices $v_{j-1}$, $v_{j+1}$ lie in $G_{1}$ and the other lie in $G_{2}$. Also, $v_{i}$ has two adjacent vertices, say $v_{i-1}\in V(G_{1})$ and $v_{i+1}\in V(G_{2})$. Since $G_{1}$ and $G_{2}$ are isomorphic, the $v_{i-1},v_{j-1}$-geodesic and the $v_{i+1},v_{j+1}$-geodesic are isomorphic, which in turn implies that $d(v_{i},v_{j-1})=d(v_{i},v_{j+1})$. This is not possible as $P$ is a geodesic. ∎
As the butterfly network is symmetrical with respect to level 00, it is symmetrical with respect to level $r$, cf. [7, 13, 16]. Using the logic of the proof of Lemma 2.3, one can prove the following.
Corollary 2.4.
If both end vertices of a geodesic $P$ of ${\rm BF}(r)$ are either at level 00 or at level $r$, then $P$ is maximal.
Lemma 2.5.
@@
Proof.
Suppose a geodesic $P$ contains four vertices $a,b,c,d$ of degree $2$. Then all the vertices $a,b,c,d$ are at level 00 or at level $r$. By Lemma 2.3, three of these vertices can not be at the same level. Assume without loss of generality that $a$ and $b$ are at level 00 and $c$ and $d$ are at level $r$. Let $P(a,b)$ be a subpath of $P$ between $a$ and $b$, and $P(c,d)$ the subpath of $P$ between $c$ and $d$. By Corollary 2.4, $P(a,b)$ and $P(c,d)$ are maximal geodesics, a contradiction. ∎
Corollary 2.6.
@@ 3.1.1 Revisiting properties of ${\rm BF}(r)$
In this section, we use the following notations. Let $U$ and $W$ denote the sets of vertices at level 00 and level $r$ in ${\rm BF}(r)$, respectively. Further, let $U=U^{b}\bigcup U^{r}$, where $U^{b}=\{u^{b}_{1},u^{b}_{2},\dots,u^{b}_{2^{r-1}}\}$ and $U^{r}=\{u^{r}_{1},u^{r}_{2},\dots,u^{r}_{2^{r-1}}\}$. Similarly, $W=W^{b}\bigcup W^{r}$ where $W^{b}=\{w^{b}_{1},w^{b}_{2},\dots,w^{b}_{2^{r-1}}\}$ and $W^{r}=\{w^{r}_{1},w^{r}_{2},\dots,w^{r}_{2^{r-1}}\}$, see Fig. 2(a).
@@
Proof.
${\rm BF}(r)-W$ consists of two components both isomorphic to ${\rm BF}(r-1)$, cf. [7, 6]. As these components are furthermore convex in ${\rm BF}(r)$, we get that if either $u^{b}_{i},u^{b}_{j}\in U^{b}$, a maximal $u^{b}_{i},u^{b}_{j}$-geodesic does not intersect $W$. Analogously the other assertions hold. The assertion (3) follows from the fact that when $u\in U$ and $w\in W$, a $u,v$-geodesic is unique. ∎
@@
$M_{U,W}({\rm BF}(r))$ is thus the set of all maximal $x,y$-geodesic, where either $x,y\in U$ or $x,y\in W$. In (i)-(iv) of Observation 3.2, given a pair of vertices $x,y\in U$ or $x,y\in W$, there are more than one maximal geodesics between $x$ and $y$ in $M_{U,W}({\rm BF}(r))$. Now we define $M^{\prime}_{U,W}({\rm BF}(r))\subset M_{U,W}({\rm BF}(r))$ as follows. First, $M^{\prime}_{U,W}({\rm BF}(r))$ contains all the geodesics from (v) and (vi) of Observation 3.2. Second, for each pair of vertices $u^{b}_{i},u^{b}_{j}\in U^{b}$ from (i), $u^{r}_{i},u^{r}_{j}\in U^{r}$ from (ii), $w^{b}_{i},w^{b}_{j}\in W^{b}$ from (iii), and $w^{r}_{i},w^{r}_{j}\in W^{r}$ from (iv), select an arbitrary but fixed geodesic between them and add it to $M^{\prime}_{U,W}({\rm BF}(r))$. In this way the set $M^{\prime}_{U,W}({\rm BF}(r))$ is defined.
@@
Note that for each pair $u_{i},u_{j}$, for each pair $w_{i},w_{j}$, for each triple $u^{b},u^{r},w$, and for each triple $w^{b},w^{r},u$, the set $M^{\prime}_{U,W}({\rm BF}(r))$ contains a unique corresponding geodesic. Note also that $M^{\prime}_{U,W}({\rm BF}(r))\subset M_{U,W}({\rm BF}(r))\subset M({\rm BF}(r))$.
@@ 3.1.2 Estimating a lower bound for ${\rm gcover}({\rm BF}(r))$
By Lemma 2.1, we get ${\rm gcover}(G)\geq{\rm gcover}[M(G),U\cup W,G]$. By Observation 3.2 and the definition of $M^{\prime}_{U,W}(G)$, ${\rm gcover}[M_{U,W}(G),U\cup W,G]={\rm gcover}[M^{\prime}_{U,W}(G),U\cup W,G]$. Next we prove that ${\rm gcover}[M(G),U\cup W,G]={\rm gcover}[M_{U,W}(G),U\cup W,G]$.
Since $M_{U,W}(G)\subseteq M(G)$, by Lemma 2.1, we get ${\rm gcover}[M(G),U\cup W,G]\leq{\rm gcover}[M_{U,W}(G),U\cup W,G]$. Now it is enough to prove that ${\rm gcover}[M(G),U\cup W,G]\geq{\rm gcover}[M_{U,W}(G),U\cup W,G]$. By Lemma 2.5, a geodesic covers at most three vertices of $U\cup W$. If $P$ is a member of $M(G)$ such that $P$ covers three vertices of $U\cup W$ in ${\rm BF}(r)$, then by Corollary 2.6, $P\in M_{U,W}(G)$. On the other hand, if $P$ is a member of $M(G)$ covering two vertices $v_{1}$ and $v_{2}$ of $U\cup W$, then by Observation 3.2 there exists a geodesic $Q$ of $M_{U,W}(G)$ such that $Q$ covers both vertices $v_{1}$ and $v_{2}$. Hence, ${\rm gcover}[M(G),U\cup W,G]\geq{\rm gcover}[M_{U,W}(G),U\cup W,G]$. ∎
@@
$\begin{split}{\rm gcover}(M^{\prime}_{U,W}(G),U\cup W,G)&\geq{\rm gcover}(M(G^{\prime\prime}),X\cup Y,G^{\prime\prime})\\ &={\rm gcover}(M(G^{\prime}),X\cup Y,G^{\prime})\\ &={\rm gcover}(G^{\prime})\\ &\geq\lceil(2/3)2^{r}\rceil\,.\\ \end{split}$
@@
Since $G^{\prime}(U,V,E^{\prime})$ is a complete bipartite graph $K_{2^{r},2^{r}}$, by Lemma 2.1 and Proposition 2.2, we infer that ${\rm gcover}(M(G^{\prime}),X\cup Y,G^{\prime})={\rm gcover}(G^{\prime})\geq\lceil(2/3)2^{r}\rceil$. ∎
@@ 3.2 An upper bound for the geodesic cover number of butterfly networks
In this section, our aim is to construct a geodesic cover of cardinality $\lceil(2/3)2^{r}\rceil$ for ${\rm BF}(r)$. In ${\rm BF}(r)$, there are $2^{r}$ rows and $r+1$ levels. The set of vertices at level 00 is $U=\{u_{1},\ldots,u_{2^{r}}\}$, the set of vertices at level $r$ in $W=\{w_{1},\ldots,w_{2^{r}}\}$. In this section, the order of vertices in $U$ and $W$ are with respect to normal representation of ${\rm BF}(r)$. (In the previous section, the order was with respect to diamond representation.) Refer to Fig. 5. The set $U$ is further partitioned into $A$ and $B$, and $W$ is partitioned into $C$ and $D$, see Fig. 5. These sets are formally defined as follows:
$A$ =$=\{[0,1],[0,2],\ldots,[0,2^{r-1}]\},$
$B$ =$=\{[0,2^{r-1}+1],[0,2^{r-1}+2],\ldots,[0,2^{r}]\},$
$C$ =$=\{[r,1],[r,2],\ldots,[r,2^{r-1}]\},$
$D$ =$=\{[r,2^{r-1}+1],[r,2^{r-1}+2],\ldots,[r,2^{r}]\}.$
@@
Property 3.6.
If a vertex $v$ is at level 00 ((((level $r)$ and vertices $u,w$ at level $r$ ((((level 0)0)0 ) are in opposite colors, then there exists a unique diametral $P_{v}(u,w)$ in ${\rm BF}(r)$.
Proof.
The structural details of two different representations of ${\rm BF}(r)$ which are illustrated in Fig. 1 are explained in [10]. By Lemma 3.1, a geodesic $P(x,y)$ between a vertex $x$ at level 00 and a vertex $y$ at level $r$ is unique in ${\rm BF}(r)$ and the length of $P(x,y)$ is $r$. From the diamond representation of ${\rm BF}(r)$ in Fig. 1 (b), whenever the vertex $v$ at level 00 ((((level $r)$ is in any color and vertices $u,w$ at level $r$ ((((level 0)0)0 ) are in opposite colors, there exists a diametral $P_{v}(u,w)$ of ${\rm BF}(r)$ between $u$ and $w$ passing through $v$. Since $P(u,v)$ and $P(v,w)$ are unique, $P_{v}(u,w)$ is also unique. ∎
Figure 5: In Stage $1$, geodesics $P_{u_{1}}(w_{1},w_{9})$, $P_{u_{2}}(w_{2},w_{10})$ and $P_{u_{16}}(w_{16},w_{8})$, $P_{u_{15}}(w_{15},w_{7})$ are constructed (pink). In Stage $2$, geodesics $P_{w_{3}}(u_{3},u_{4})$, $P_{w_{5}}(u_{5},u_{6})$ and $P_{w_{14}}(u_{14},u_{13})$, $P_{w_{12}}(u_{12},u_{11})$ are constructed (green). The vertices which are not covered in previous stages are circled. They are then covered in Stage $3$.
Thus, by Property 3.6, in order to construct a diametral path $P_{v}(u,w)$ in ${\rm BF}(r)$, it is enough to identify the middle vertex $v$ at level 00 (level $r$) in any color and end vertices $\{u,w\}$ at level $r$ (level 00) in opposite colors.
@@
    $U^{r}=\{u^{r}_{i}:\ i\in[2^{r-3}]\}$ the red vertices of $A^{\prime}$ and $B^{\prime}$.
@@
    $U^{b}=\{u^{b}_{i}:\ i\in[2^{r-3}]\}$ the blue vertices of $A^{\prime}$ and $B^{\prime}$.
@@
    $W^{b}=\{w^{b}_{i}:\ i\in[2^{r-3}]\}$ the vertices of $D^{\prime}$.
(The sets $W^{r}$ and $W^{b}$ are thus obtained by renaming the vertices of $C^{\prime}$ and $D^{\prime}$.) The next step in Stage 3 is to partition the vertices of $U^{r}$, $U^{b}$, $W^{r}$, and $W^{b}$ into four subsets. For a fixed $r\geq 5$, let $\ell=\lfloor\frac{2^{r-3}}{3}\rfloor$. Then $12^{r-3}=3\cdot\ell+$ 1 or $22^{r-3}=3\cdot\ell+$ 2.
Case 1: $12^{r-3}=3\cdot\ell+$ 1.
@@
$U^{r}$ $U^{r}_{1}$ $u^{r}_{1},$ $u^{r}_{2},$ $\ldots$,… , $u^{r}_{\ell}$ $U^{r}_{2}$ $u^{r}_{\ell+1},$ $u^{r}_{\ell+2}$, $\ldots$,… , $u^{r}_{2\ell}$ $U^{r}_{3}$ $u^{r}_{2\ell+1},$ $u^{r}_{2\ell+2}$, $\ldots$,… , $u^{r}_{3\ell}$ $U^{r}_{4}$ $u^{r}_{x}$ $W^{r}$ $W^{r}_{1}$ $w^{r}_{1},$ $w^{r}_{2}$, $\ldots$, $w^{r}_{\ell}$ $W^{r}_{2}$ $w^{r}_{\ell+1},$ $w^{r}_{\ell+2}$, $\ldots$, $w^{r}_{2\ell}$ $W^{r}_{3}$ $w^{r}_{2\ell+1},$ $w^{r}_{2\ell+2}$, $\ldots$, $w^{r}_{3\ell}$ $W^{r}_{4}$ $w^{r}_{x}$
@@
Figure 6: Here $U=\{u^{r}_{1},u^{r}_{2},u^{r}_{3},u^{b}_{1},u^{b}_{2},u^{b}_{3}\}$ and $W=\{w^{r}_{1},w^{r}_{2},w^{r}_{3},w^{b}_{1},w^{b}_{2},w^{b}_{3}\}$, where $u^{r}_{1},u^{r}_{2},u^{r}_{3},w^{r}_{1},w^{r}_{2},w^{r}_{3}$ are red vertices, and $u^{b}_{1},u^{b}_{2},u^{b}_{3},w^{b}_{1},w^{b}_{2},w^{b}_{3}$ are blue vertices. How to cover the vertices by geodesics of length $3$ with end vertices in opposite colors?
@@
In Case $1$ of Stage $3$ we have thus constructed $24\ell+$ 2 $=\lceil\frac{2^{r-1}}{3}\rceil$ geodesics covering all the vertices of $U^{r}\cup U^{b}\cup W^{r}\cup W^{b}=A^{\prime}\cup B^{\prime}\cup C^{\prime}\cup D^{\prime}$.
Case 2: $22^{r-3}=3\cdot\ell+$ 2.

### v3

@@ Abstract
volume: 194issue: 3
The geodesic cover problem for butterfly networks
@@ 2 Preliminaries
    If $S^{\prime}\subseteq S^{\prime\prime}\subseteq V(G)$ and $Y\subseteq Z(G)$, then ${\rm gcover}(Y,S^{\prime\prime},G)$ $\geq$ ${\rm gcover}(Y,S^{\prime},G)$.
@@
    ${\rm gcover}(G)$ == ${\rm gcover}(M(G),G)$.
Proof 2.2
Assertions (i) and (ii) are straightforward, hence we consider only (iii). Since $M(G)\subseteq Z(G)$, (ii) implies
${\rm gcover}(G)={\rm gcover}(Z(G),V,G)\leq{\rm gcover}(M(G),V,G)={\rm gcover}(M(G),G).$
Since each geodesic is a subpath of some maximal geodesic, for each geodesic cover $S$ of $Z(G)$, there exists a geodesic cover $S^{\prime}$ of $M(G)$ such that $|S|=|S^{\prime}|$. Therefore,
${\rm gcover}(Z(G),V,G)\geq{\rm gcover}(M(G),V,G)$
and consecutively
${\rm gcover}(G)={\rm gcover}(Z(G),V,G)\geq{\rm gcover}(M(G),V,G)={\rm gcover}(M(G),G).$
Proposition 2.3
@@
Proof 2.4
Clearly, each maximal geodesic of $K_{r,r}$ is a (diametral) path of length $2$. Therefore, ${\rm gcover}(K_{r,r})\geq\lceil(2/3)r\rceil$. On the other hand, it is a simple exercise to construct a geodesic cover of cardinality $\lceil(2/3)r\rceil$.
Butterfly is considered as one of the best parallel architectures [6, 7, 15]. For $r\geq 3$, the $r$-dimensional butterfly network $BF(r)$ has vertices $[j,s]$, where $s\in\{0,1\}^{r}$ and $j\in\{0,1,\ldots,r\}$. The vertices $[j,s]$ and $j^{\prime},s^{\prime}]$ are adjacent if $j^{\prime}=j+1$, and either $s=s^{\prime}$ or $s$ and $s^{\prime}$ differ precisely in the $j^{\rm th}$ bit. $BF(r)$ has $(r+1)2^{r}$ vertices and $r2^{r+1}$ edges. A vertex $[j,s]$ is at level $j$ and row $s$. There are two standard graphical representations for ${\rm BF}(r)$, normal representation and diamond representation, see Fig. 1.
@@
Lemma 2.5
A geodesic of ${\rm BF}(r)$ contains at most two vertices of level 0 and at most two vertices of level $r$. Moreover, if a geodesic contains two vertices of level 0, then they are the ends of the geodesic (and similarly for level r).
Proof 2.6
Let us assume that there exists a geodesic $P$ which contains more than two vertices of level 0, say $v_{i},v_{j}$, and $v_{k}$. See Fig. 2(b). Then one of these three vertices must be an internal vertex of $P$, say $v_{j}$. The deletion of the vertices at level 0 from ${\rm BF}(r)$ disconnects ${\rm BF}(r)$ into two vertex disjoint components $G_{1}$ and $G_{2}$, where both $G_{1}$ and $G_{2}$ are isomorphic to ${\rm BF}(r-1)$, cf. [7, 13, 16]. Since $v_{j}$ is an internal vertex of $P$ and of degree $2$, its neighbors $v_{j-1}$ and $v_{j+1}$ also lie in $P$. Moreover, one of the adjacent vertices $v_{j-1}$, $v_{j+1}$ lie in $G_{1}$ and the other lie in $G_{2}$. Also, $v_{i}$ has two adjacent vertices, say $v_{i-1}\in V(G_{1})$ and $v_{i+1}\in V(G_{2})$. Since $G_{1}$ and $G_{2}$ are isomorphic, the $v_{i-1},v_{j-1}$-geodesic and the $v_{i+1},v_{j+1}$-geodesic are isomorphic, which in turn implies that $d(v_{i},v_{j-1})=d(v_{i},v_{j+1})$. This is not possible as $P$ is a geodesic.
As the butterfly network is symmetrical with respect to level 0, it is symmetrical with respect to level $r$, cf. [7, 13, 16]. Using the logic of the proof of Lemma 2.5, one can prove the following.
Corollary 2.7
If both end vertices of a geodesic $P$ of ${\rm BF}(r)$ are either at level 0 or at level $r$, then $P$ is maximal.
Lemma 2.8
@@
Proof 2.9
Suppose a geodesic $P$ contains four vertices $a,b,c,d$ of degree $2$. Then all the vertices $a,b,c,d$ are at level 0 or at level $r$. By Lemma 2.5, three of these vertices can not be at the same level. Assume without loss of generality that $a$ and $b$ are at level 0 and $c$ and $d$ are at level $r$. Let $P(a,b)$ be a subpath of $P$ between $a$ and $b$, and $P(c,d)$ the subpath of $P$ between $c$ and $d$. By Corollary 2.7, $P(a,b)$ and $P(c,d)$ are maximal geodesics, a contradiction.
Corollary 2.10
@@ 3.1.1 Revisiting properties of ${\rm BF}(r)$
In this section, we use the following notations. Let $U$ and $W$ denote the sets of vertices at level 0 and level $r$ in ${\rm BF}(r)$, respectively. Further, let $U=U^{b}\bigcup U^{r}$, where $U^{b}=\{u^{b}_{1},u^{b}_{2},\dots,u^{b}_{2^{r-1}}\}$ and $U^{r}=\{u^{r}_{1},u^{r}_{2},\dots,u^{r}_{2^{r-1}}\}$. Similarly, $W=W^{b}\bigcup W^{r}$ where $W^{b}=\{w^{b}_{1},w^{b}_{2},\dots,w^{b}_{2^{r-1}}\}$ and $W^{r}=\{w^{r}_{1},w^{r}_{2},\dots,w^{r}_{2^{r-1}}\}$, see Fig. 2(a).
@@
Proof 3.2
${\rm BF}(r)-W$ consists of two components both isomorphic to ${\rm BF}(r-1)$, cf. [7, 6]. As these components are furthermore convex in ${\rm BF}(r)$, we get that if either $u^{b}_{i},u^{b}_{j}\in U^{b}$, a maximal $u^{b}_{i},u^{b}_{j}$-geodesic does not intersect $W$. Analogously the other assertions hold. The assertion (3) follows from the fact that when $u\in U$ and $w\in W$, a $u,v$-geodesic is unique.
@@
$M_{U,W}({\rm BF}(r))$ is thus the set of all maximal $x,y$-geodesic, where either $x,y\in U$ or $x,y\in W$. In (i)-(iv) of Observation 3.1.1, given a pair of vertices $x,y\in U$ or $x,y\in W$, there are more than one maximal geodesics between $x$ and $y$ in $M_{U,W}({\rm BF}(r))$. Now we define $M^{\prime}_{U,W}({\rm BF}(r))\subset M_{U,W}({\rm BF}(r))$ as follows. First, $M^{\prime}_{U,W}({\rm BF}(r))$ contains all the geodesics from (v) and (vi) of Observation 3.1.1. Second, for each pair of vertices $u^{b}_{i},u^{b}_{j}\in U^{b}$ from (i), $u^{r}_{i},u^{r}_{j}\in U^{r}$ from (ii), $w^{b}_{i},w^{b}_{j}\in W^{b}$ from (iii), and $w^{r}_{i},w^{r}_{j}\in W^{r}$ from (iv), select an arbitrary but fixed geodesic between them and add it to $M^{\prime}_{U,W}({\rm BF}(r))$. In this way the set $M^{\prime}_{U,W}({\rm BF}(r))$ is defined.
@@
Note that for each pair $u_{i},u_{j}$, for each pair $w_{i},w_{j}$, for each triple $u^{b},u^{r},w$, and for each triple $w^{b},w^{r},u$, the set $M^{\prime}_{U,W}({\rm BF}(r))$ contains a unique corresponding geodesic. Note also that
$M^{\prime}_{U,W}({\rm BF}(r))\subset M_{U,W}({\rm BF}(r))\subset M({\rm BF}(r)).$
@@ 3.1.2 Estimating a lower bound for ${\rm gcover}({\rm BF}(r))$
By Lemma 2.1, we get ${\rm gcover}(G)\geq{\rm gcover}[M(G),U\cup W,G]$. By Observation 3.1.1 and the definition of $M^{\prime}_{U,W}(G)$, ${\rm gcover}[M_{U,W}(G),U\cup W,G]={\rm gcover}[M^{\prime}_{U,W}(G),U\cup W,G]$. Next we prove that ${\rm gcover}[M(G),U\cup W,G]={\rm gcover}[M_{U,W}(G),U\cup W,G]$.
Since $M_{U,W}(G)\subseteq M(G)$, by Lemma 2.1, we get
${\rm gcover}[M(G),U\cup W,G]\leq{\rm gcover}[M_{U,W}(G),U\cup W,G].$
Now it is enough to prove that
${\rm gcover}[M(G),U\cup W,G]\geq{\rm gcover}[M_{U,W}(G),U\cup W,G].$
By Lemma 2.8, a geodesic covers at most three vertices of $U\cup W$. If $P$ is a member of $M(G)$ such that $P$ covers three vertices of $U\cup W$ in ${\rm BF}(r)$, then by Corollary 2.10, $P\in M_{U,W}(G)$. On the other hand, if $P$ is a member of $M(G)$ covering two vertices $v_{1}$ and $v_{2}$ of $U\cup W$, then by Observation 3.1.1 there exists a geodesic $Q$ of $M_{U,W}(G)$ such that $Q$ covers both vertices $v_{1}$ and $v_{2}$. Hence, ${\rm gcover}[M(G),U\cup W,G]\geq{\rm gcover}[M_{U,W}(G),U\cup W,G]$.
@@
[…]
@@
Since $G^{\prime}(U,V,E^{\prime})$ is a complete bipartite graph $K_{2^{r},2^{r}}$, by Lemma 2.1 and Proposition 2.3, we infer that ${\rm gcover}(M(G^{\prime}),X\cup Y,G^{\prime})={\rm gcover}(G^{\prime})\geq\lceil(2/3)2^{r}\rceil$.
@@ 3.2 An upper bound for the geodesic cover number of butterfly networks
In this section, our aim is to construct a geodesic cover of cardinality $\lceil(2/3)2^{r}\rceil$ for ${\rm BF}(r)$. In ${\rm BF}(r)$, there are $2^{r}$ rows and $r+1$ levels. The set of vertices at level 0 is $U=\{u_{1},\ldots,u_{2^{r}}\}$, the set of vertices at level $r$ in $W=\{w_{1},\ldots,w_{2^{r}}\}$. In this section, the order of vertices in $U$ and $W$ are with respect to normal representation of ${\rm BF}(r)$. (In the previous section, the order was with respect to diamond representation.) Refer to Fig. 5. The set $U$ is further partitioned into $A$ and $B$, and $W$ is partitioned into $C$ and $D$, see Fig. 5. These sets are formally defined as follows:
$A$ $=\{[0,1],[0,2],\ldots,[0,2^{r-1}]\},$
$B$ $=\{[0,2^{r-1}+1],[0,2^{r-1}+2],\ldots,[0,2^{r}]\},$
$C$ $=\{[r,1],[r,2],\ldots,[r,2^{r-1}]\},$
$D$ $=\{[r,2^{r-1}+1],[r,2^{r-1}+2],\ldots,[r,2^{r}]\}.$
@@
Property 3.7
If a vertex $v$ is at level 0 ((level $r)$ and vertices $u,w$ at level $r$ ((level $0)$ are in opposite colors, then there exists a unique diametral $P_{v}(u,w)$ in ${\rm BF}(r)$.
Proof 3.8
The structural details of two different representations of ${\rm BF}(r)$ which are illustrated in Fig. 1 are explained in [10]. By Lemma 3.1, a geodesic $P(x,y)$ between a vertex $x$ at level 0 and a vertex $y$ at level $r$ is unique in ${\rm BF}(r)$ and the length of $P(x,y)$ is $r$. From the diamond representation of ${\rm BF}(r)$ in Fig. 1 (b), whenever the vertex $v$ at level 0 ((level $r)$ is in any color and vertices $u,w$ at level $r$ ((level $0)$ are in opposite colors, there exists a diametral $P_{v}(u,w)$ of ${\rm BF}(r)$ between $u$ and $w$ passing through $v$. Since $P(u,v)$ and $P(v,w)$ are unique, $P_{v}(u,w)$ is also unique.
Figure 5: In Stage 11, geodesics $P_{u_{1}}(w_{1},w_{9})$, $P_{u_{2}}(w_{2},w_{10})$ and $P_{u_{16}}(w_{16},w_{8})$, $P_{u_{15}}(w_{15},w_{7})$ are constructed (pink). In Stage 22, geodesics $P_{w_{3}}(u_{3},u_{4})$, $P_{w_{5}}(u_{5},u_{6})$ and $P_{w_{14}}(u_{14},u_{13})$, $P_{w_{12}}(u_{12},u_{11})$ are constructed (green). The vertices which are not covered in previous stages are circled. They are then covered in Stage 33.
Thus, by Property 3.2, in order to construct a diametral path $P_{v}(u,w)$ in ${\rm BF}(r)$, it is enough to identify the middle vertex $v$ at level 0 (level $r$) in any color and end vertices $\{u,w\}$ at level $r$ (level 0) in opposite colors.
@@
    $U^{r}=\{u^{r}_{i}:\ i\in[2^{r-3}]\}$ - the red vertices of $A^{\prime}$ and $B^{\prime}$.
@@
    $U^{b}=\{u^{b}_{i}:\ i\in[2^{r-3}]\}$ - the blue vertices of $A^{\prime}$ and $B^{\prime}$.
@@
    $W^{b}=\{w^{b}_{i}:\ i\in[2^{r-3}]\}$ - the vertices of $D^{\prime}$.
(The sets $W^{r}$ and $W^{b}$ are thus obtained by renaming the vertices of $C^{\prime}$ and $D^{\prime}$.) The next step in Stage 3 is to partition the vertices of $U^{r}$, $U^{b}$, $W^{r}$, and $W^{b}$ into four subsets. For a fixed $r\geq 5$, let $\ell=\lfloor\frac{2^{r-3}}{3}\rfloor$. Then $2^{r-3}=3\cdot\ell+1$ or $2^{r-3}=3\cdot\ell+2$.
Case 1: $2^{r-3}=3\cdot\ell+1$.
@@
$U^{r}$ $U^{r}_{1}$ $u^{r}_{1},$ $u^{r}_{2},$ $\ldots,$ $u^{r}_{\ell}$ $U^{r}_{2}$ $u^{r}_{\ell+1},$ $u^{r}_{\ell+2}$, $\ldots,$ $u^{r}_{2\ell}$ $U^{r}_{3}$ $u^{r}_{2\ell+1},$ $u^{r}_{2\ell+2}$, $\ldots,$ $u^{r}_{3\ell}$ $U^{r}_{4}$ $u^{r}_{x}$ $W^{r}$ $W^{r}_{1}$ $w^{r}_{1},$ $w^{r}_{2}$, $\ldots$, $w^{r}_{\ell}$ $W^{r}_{2}$ $w^{r}_{\ell+1},$ $w^{r}_{\ell+2}$, $\ldots$, $w^{r}_{2\ell}$ $W^{r}_{3}$ $w^{r}_{2\ell+1},$ $w^{r}_{2\ell+2}$, $\ldots$, $w^{r}_{3\ell}$ $W^{r}_{4}$ $w^{r}_{x}$
@@
Figure 6: Here $U=\{u^{r}_{1},u^{r}_{2},u^{r}_{3},u^{b}_{1},u^{b}_{2},u^{b}_{3}\}$ and $W=\{w^{r}_{1},w^{r}_{2},w^{r}_{3},w^{b}_{1},w^{b}_{2},w^{b}_{3}\}$, where $u^{r}_{1},u^{r}_{2},u^{r}_{3},w^{r}_{1},w^{r}_{2},w^{r}_{3}$ are red vertices, and $u^{b}_{1},u^{b}_{2},u^{b}_{3},w^{b}_{1},w^{b}_{2},w^{b}_{3}$ are blue vertices. How to cover the vertices by geodesics of length 33 with end vertices in opposite colors?
@@
In Case $1$ of Stage $3$ we have thus constructed $4\ell+2$ $\lceil\frac{2^{r-1}}{3}\rceil$ geodesics covering all the vertices of $U^{r}\cup U^{b}\cup W^{r}\cup W^{b}=A^{\prime}\cup B^{\prime}\cup C^{\prime}\cup D^{\prime}$.
Case 2: $2^{r-3}=3\cdot\ell+2$.